Fermi National Laboratory


Accelerator Update

Friday November 21
The day shift began with the Antiproton Source (Pbar) stacking, with experts conducting studies in the Tevatron (TeV) and Recycler (RR), and with MiniBooNE taking beam.

At 3 PM, Pbar studiers switched from stacking to reverse proton studies.

TeV experts tried to load a 36x0 study store at 7:46 PM, but there were problems with the first injection.

At 7 PM, a Linac quadrupole power supply tripped off for the first of many times. Operations called in a technician who had the supply operating by 10:30 PM.

At 8 PM, problems developed with at breaker panel at the Pbar AP0 service building that tripped off devices in AP2 and AP3. Operations called for the Duty Electrician.

MiniBooNE had stopped receiving beam when the Linac quadrupole power supply problem appeared. Beam resumed at 10:30 PM.

Operators determined that a problem with the TeV sequencer was disrupting the study store. An expert suggested a fix and Operations put the TeV through a dry squeeze, but by 11:30 PM they still hadn't been able to inject beam.

Saturday November 22
The midnight shift began with experts conducting studies in TeV and Pbar, and with MiniBooNE taking beam.

Operators discovered the problem that was holding off TeV injections at 1 AM.

At 2:00 AM, Pbar experts returned Pbar to stacking mode, but then began Debuncher aperture scans. They resumed stacking at 4 AM.

TeV experts and operators established a 36x0 study store at 4:30 AM.

At 5:34 AM, the TeV suffered a sector A4 quench. Operators made the Collider Detector Facility (CDF) and D-zero (D0) ready for access at 6:30 AM while the cryo systems recovered from the quench.

Operations put the TeV through a dry squeeze at 9:21 AM.

A Main Injector RF station (MIRF3) tripped off due to a bad modulator fan. At 9:22 AM operators called in an expert to make repairs. He repaired the station and had Pbar stacking by 10:43 AM.

At 11:44 AM, Operations put the TeV into shot setup to load a 36x4 study store. Unfortunately the TeV quenched at sector E4 just before the antiprotons were to be loaded.

Pbar resumed stacking at 2:13 PM.

A new TeV shot setup began 4 PM. Three of the four antiproton injections made it into the TeV. The initial luminosity was negligible.

At 7:16 PM, an H- Source vacuum pump tripped off due to overheating. Operators opened the back of the equipment rack and that cooled off the pump enough to resume operating.

At 7:30 PM, RR experts began a study period and Pbar experts began Debuncher reverse proton studies.

At 10 PM, a Booster RF station (BRF14) tripped off and then at 10:30 PM BRF01 and a Linac RF station (LRF5) tripped off. They all reset okay.

Pbar studies ended and they resumed stacking at 11:42 PM.

Sunday November 23
The midnight shift began with Pbar stacking, with a TeV study store, and with MiniBooNE taking beam.

Operations terminated the TeV study store at 12:14 AM.

At 4:30 AM, a Linac RF station (KRF1) tripped off. It reset okay.

KRF5 tripped off at 9:30 AM for the first of three times over a forty-five minute period. It reset okay each time.

At 1 PM, a Pbar expert began a stacking study.

At 2:22 PM, Operations noticed the H- Source dome sparking. They accessed and discovered a leak from all the rain. After consulting an expert, operators switched Preacc operation to the I- Source.

At 9:41 PM, TeV experts were completing their final chromaticity check, Pbar was completing their final tune, and MI was ready and waiting.

Operations and experts put the TeV into shot setup. At 11 PM, they established a study store with an initial luminosity of 0.9E31. Pbar resumed stacking.

Monday November 24
The midnight shift began with Operations monitoring the study store, with Pbar stacking, and with MiniBooNE taking beam.

At 12:10 AM, a Booster RF station tripped off due to no LCW flow. Operators investigated and found that there was flow. They bypassed the alarm and continued running. The Booster will have to be accessed at the next opportunity to investigate the cavity's flow-switch.

A Linac RF station tripped off at 2:16 AM due to its water skid tripping off. Operators called in a tech and he had the station back online by 4:47 AM. Stacking resumed.

Two Booster RF station (BRF13 & 14) tripped off due to a vacuum burst at 6 AM. They reset okay.

At 6:34 AM, Operations received a bunch of alarms for the Meson area and CHL called up saying they'd been hit by a power glitch. Operations called the Duty Electrician. The high voltage expert reported that feeder 30 faulted. Unfortunately, feeder 31 was down for repairs so there's no backup. CHL was recovering.

Today's Plans
The study store will be terminated early this morning. Then accesses into various parts of the accelerator complex will be allowed for necessary repairs. After the accesses Operations will establish a store. Technicians will finish their work on feeder 31 this evening. The feeder 30 fault hasn't been found.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 11/24/2003   email Fermilab